# Mitigating Cache Noise in Test-Time Adaptation for Large Vision-Language Models

### v1

@@ 4 Experiment
We follow previous work[13] to evaluate our method on two benchmarking scenarios: cross-dataset generalization and robustness to natural distribution shifts. (1) For cross-dataset generalization tasks, we conduct comprehensive assessments across 10 diverse recognition datasets, including FGVC Aircraft[18], Caltech101[6], Stanford Cars[15], DTD[4], EuroSAT[11], Flowers102[19], Food101[3], Oxford Pets[20], SUN397[27], and UCF101[25]. These datasets provide a comprehensive benchmark for evaluating the adaptability and generalization ability of methods across different datasets. (2) For evaluating robustness to natural distribution shifts, we assess the performance of our method using the ImageNet[5] dataset alongside its variant out-of-distribution datasets, including ImageNet-A[12], ImageNet-V2[23]. This evaluation measures our method’s robustness in the presence of different distribution shifts.It should be noted that for datasets with excessively large test sets, we only extract the 2000 samples from the test set, similar to DiffTPT[7].
Table 1: Performance comparisons on robustness to natural distribution shifts. The best results are highlighted in bold. The asterisk (*) indicates that our method is evaluated on a subset of the test set.
@@ 4.3 Comparisons with State-of-the-art
Robustness to Natural Distribution Shifts. Our approach consistently demonstrates strong generalization performance on downstream tasks with natural distribution shifts. As shown in Table 1, our method achieved the highest average results across datasets with significant distribution differences, each evaluated using different backbone networks. Specifically, we achieved the best performance on ImageNet, with improvements of 3.91 and 5.5 percentage points. It is noteworthy that the CoOp results were obtained using few-shot learning. Our method not only surpasses TDA but in certain cases also outperforms few-shot methods, showcasing the superiority of our approach. Although there are instances where our performance is not as strong as DiffTPT, we speculate this may be due to the lack of extensive data augmentation. Our method can be integrated with data augmentation techniques, but this was not the primary focus of our study.
Table 2: Performance comparisons on cross-datesets generalization. The best results are highlighted in bold. The asterisk (*) indicates that our method is evaluated on a subset of the test set.
@@
Ours 26.58 93.57 66.89 53.78 59.81 75.94 84.81* 91.20 68.36* 70.31 69.12
Cross-Datasets Generalization. In Table 2, we further evaluate the generalization capability of our proposed method against other state-of-the-art methods across 10 fine-grained recognition datasets. Due to the substantial distribution differences among these datasets, performance can vary considerably. Despite this, our method achieves average improvements of 1.39 and 1.59 over the current best-performing methods on two distinct backbones, surpassing them on 6 out of the 10 datasets.
Notably, we observe a marked performance gain on the DTD dataset, likely because its strong emphasis on textures makes the underlying data distribution crucial for texture description. Conversely, our results on Food101 lag behind the current SOTA, which may be attributed to the high inter-class similarity that challenges GDA in accurately modeling class distributions. Overall, these findings underscore the remarkable robustness and adaptability of our approach when transferring to new domains at test time—a critical factor for real-world applications.
@@ Appendix A Appendix
  • Preliminaries and Methodological Differences: A discussion of a foundational concept is included, along with an explanation of the key distinctions between our approach and TDA[13].
@@ A.2 Preliminaries and Methodological Differences
Differences from TDA Although our approach shares many visual similarities with TDA[13] methods, it is fundamentally different in methodology. Our Negative Cache Prototypes are entirely derived from the Positive Cache and are constructed following the principles outlined in the main text. This means we focus exclusively on high-confidence images. Building on this foundation, we further learn a bi-modal residual to better handle each test sample. Additionally, we incorporate Gaussian Discriminant Analysis during inference, which effectively mitigates the impact of noisy labels. The differences between our method and those of TDA and TPT are clearly illustrated in 3

### v2

@@ 4 Experiment
We follow previous work[13] to evaluate our method on two benchmarking scenarios: cross-dataset generalization and robustness to natural distribution shifts. (1) For cross-dataset generalization tasks, we conduct comprehensive assessments across 10 diverse recognition datasets, including FGVC Aircraft[18], Caltech101[6], Stanford Cars[15], DTD[4], EuroSAT[11], Flowers102[19], Food101[3], Oxford Pets[20], SUN397[29], and UCF101[25]. These datasets provide a comprehensive benchmark for evaluating the adaptability and generalization ability of methods across different datasets. (2) For evaluating robustness to natural distribution shifts, we assess the performance of our method using the ImageNet[5] dataset alongside its variant out-of-distribution datasets, including ImageNet-A[12], ImageNet-V2[23]. This evaluation measures our method’s robustness in the presence of different distribution shifts. For datasets with excessively large test sets—Food101, ImageNet, SUN397, ImageNet-V2, and ImageNet-A—we randomly sample 2000 test instances for evaluation, following the experimental setting in DiffTPT[7].
Table 1: Performance comparisons on robustness to natural distribution shifts. The best results are highlighted in bold.
@@ 4.3 Comparisons with State-of-the-art
Robustness to Natural Distribution Shifts. Our approach consistently demonstrates strong generalization performance on downstream tasks with natural distribution shifts. As shown in Table 1, our method achieved the highest average results across datasets with significant distribution differences, each evaluated using different backbone networks. Specifically, we achieved the best performance on ImageNet, with improvements of 1.85 and 3.1 percentage points. It is noteworthy that the CoOp results were obtained using few-shot learning. Our method not only surpasses TDA and DPE but, in certain cases, also outperforms few-shot methods, showcasing the superiority of our approach. Although there are instances where our performance is not as strong as DiffTPT, we speculate this may be due to the lack of extensive data augmentation. Our method can be integrated with data augmentation techniques, but this was not the primary focus of our study.
Table 2: Performance comparisons on cross-datesets generalization. The best results are highlighted in bold.
@@
DPE [35] 28.95 94.81 67.31 54.20 55.79 75.07 86.17 91.14 70.07 70.44 69.40
Ours 26.58 93.57 66.89 56.38 59.81 75.94 85.95 91.20 68.36 70.31 69.50
Cross-Datasets Generalization. In Table 2, we further evaluate the generalization capability of our proposed method against other state-of-the-art methods across 10 fine-grained recognition datasets. Due to the substantial distribution differences among these datasets, performance can vary considerably. Despite this, our method achieves average improvements of 0.49 over the current best-performing methods on the CLIP-ResNet-50 backbone, surpassing them on 3 out of the 10 datasets. On the CLIP-ViT backbone, our method’s average accuracy is 0.1 higher than DPE, and is optimal on 4 of the 10 datasets. Notably, we significantly improved EuroSAT due to its more compact category distributions, yielding more representative prototypes. However, our performance on Food101 is slightly lower, likely due to the high inter-class similarity that poses challenges for GDA. Overall, these findings confirm our method’s robustness and adaptability when transferring to new domains at test time, making it crucial for real-world applications.
@@ Appendix A Appendix
  • Preliminaries and Methodological Differences: A discussion of a foundational concept is included, along with an explanation of the key distinctions between our approach and TDA[13],SimNL[34],GDA[28],DPE[35].
@@ A.2 Preliminaries and Methodological Differences
Differences from TDA[13] Although our approach shares many visual similarities with TDA[13] methods, it is fundamentally different in methodology. Our Negative Cache Prototypes are entirely derived from the Positive Cache and are constructed following the principles outlined in the main text. This means we focus exclusively on high-confidence images. Building on this foundation, we further learn a bi-modal residual to better handle each test sample. Additionally, we incorporate Gaussian Discriminant Analysis during inference, which effectively mitigates the impact of noisy labels. The differences between our method and those of TDA and TPT are clearly illustrated in 3
Differences from SimNL[34] SimNL proposes a simple yet effective negative learning method for few-shot scenarios by constructing a negative cache to learn complementary features, thereby adapting VLMs to downstream tasks. Specifically, SimNL constructs negative features on both the text and visual modalities and tunes the model by learning multimodal residuals.
Our work is inspired by the high-level idea of SimNL; however, there are significant differences in the implementation details and underlying motivation. First, the methodological paradigm is different: SimNL belongs to few-shot methods, whereas our approach focuses on test-time adaptation. We achieve unsupervised optimization by adjusting the residual between positive and negative prototypes and by incorporating an additional loss during testing to constrain their similarity, thereby forcing the two prototypes to diverge. Second, the strategy for constructing negative samples differs: we only construct negative samples on the visual modality, while SimNL constructs negative features on both text and visual modalities. Finally, regarding the construction of negative prototypes, SimNL obtains them by randomly sampling images from the $C-1$ classes other than the target class and averaging them, repeating this process $K$ times to form a negative cache of size $CK\times D$ (the same shape as the positive sample cache), and then fusing the residuals via broadcasting for learning, whereas our method directly computes the negative prototype from the positive prototype, resulting in a negative prototype of size $C\times D$. Moreover, our motivation for introducing negative learning is to suppress label noise, which further distinguishes our approach from that of SimNL.
Differences from GDA[28] The work in [28] is the first to introduce Gaussian Discriminant Analysis (GDA) into the adaptation of VLMs. We acknowledge that there is a degree of similarity between their approach and ours, as both employ GDA. However, there are several important differences in our respective settings, motivations, and technical details. First, [28] focuses on few-shot scenarios, extending to base-to-new class generalization and unsupervised learning, whereas our method is designed for test-time adaptation using a cache-based strategy. Second, while [28] primarily adopts GDA as a training-free solution to minimize additional parameters and computational overhead, we leverage GDA from a distributional perspective to enhance robustness to noisy labels—a notable advantage over KNN-like approaches[36]. Finally, when computing the mean and covariance for each class, we incorporate a learned residual from test-time adaptation: for each image in the cache, we add the learned residual to its representation before deriving the class mean and covariance. This residual-based adjustment further distinguishes our method from that of [28].
Difference form DPE[35] Residual learning has demonstrated great potential in the adaptation of VLMs. Taskres[32], in few-shot scenarios, achieves rapid adaptation to downstream tasks by learning only a single layer of prior-agnostic parameters on the text modality. TPS[26] extends this strategy further into test-time adaptation, adhering to the TPT[24] paradigm by continuously updating text prototypes during testing. On the other hand, DPE[35] integrates cache-based approaches with residual learning, introducing for the first time the concept of Dual Prototype Evolution in test-time adaptation. Specifically, DPE learns residuals for both textual and visual prototypes simultaneously, dynamically updating prototypes across the two modalities throughout testing. This makes test-time adaptation of vision-language models simultaneously accumulative and multimodal for the first time.
Our work is inspired by the overall framework of DPE and similarly adopts the dynamic updating mechanism for text and visual prototypes. However, our method further introduces negative prototype learning. Specifically, during test-time adaptation, we not only aim to enlarge the distinction among different textual prototypes but also explicitly reduce the similarity between positive and negative prototypes. After residual learning, we integrate the learned residuals into cached features and utilize Gaussian Discriminant Analysis for inference, thus achieving a more robust and accurate test-time adaptation performance.